OpenStudy (anonymous):

find the volume between the cylinders. R=12in. r=9in. H=20in.

OpenStudy (anonymous):

volume = volume of outer cylinder - volume of inner cylinder = πR^2H - πr^2H = πH (R^2 - r^2) = 3.14 x 20 in (12^2 - 9^2) in^2 = 62.8 (144 - 81) in^3 = 3956.4 inch^3

OpenStudy (anonymous):

where is my medal pls ::))